[figure: 9.7 cm]

VARIATION OF MAGNETIC PROPERTIES AS FUNCTION OF

COMPOZITION IN THE Gd(Fe_{1−x}Al_{x})_{11}TiC_{1−δ} AND

Gd(Fe_{1−x}Al_{x})_{10}V_{2}C_{1−δ} SERIES

Corneliu-Bazil Cizmas Transilvania University, Physics Dept., Brasov-2200, Romania

Some magnetic properties of the Gd(Fe_{1−x}Al_{x})_{11}TiC_{y} and Gd(Fe_{1−x}Al_{x})_{10}V_{2}C_{y} alloys (with x=0; 0,1; 0,2 and y=0 and 1 for x = 0,2) have been investigated using a Faraday magnetic balance in the 80-1100 K temperature range and in an external magnetic field up to 1T. The alloys with nominal compozition have been prepared by arc melting method and annealed for several days at 1000 ^{o}C. The structure and termomagnetic analyze show that the alloys are predominantly formed by 1:12 ordered phase (ThMn_{12} type structure), but low concentrations of the disordered α-Fe phase are present. From the starting compounds of the series (GdFe_{11}Ti and GdFe_{10}V_{2}) the Curie temperature is 604 K and 628 K, respectively. It is found that the Curie temperature, the saturation magnetization and the magnetic anisotropy at room temperature of the 1:12 intermetallic compounds decrease with the increase of Al content, but clearly increase by the insertion of the interstitial carbon atoms in the Gd(Fe_{0.8}Al_{0.2})_{11}TiC_{1−δ} and Gd(Fe_{0.8}Al_{0.2})_{10}V_{2}C_{1−δ} compounds. For exemple, Curie temperature decreases with Al content up to 470 K and 376 K and increase by C-atoms insertion at 504 K and 512 K respectively. The experimental data were interpreted in terms of the exchange interactions theory.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : dr. Corneliu-Bazil Cizmas

Address for correspondence : Transilvania University, Phys. Dept. Eroilor 29 Brasov 2200, ROMANIA

Email address : cizmascb@unitbv.ro